# Estimating the strength of Lorentzian distribution in non-commutative geometry by solar system tests

### v3

@@ Abstract
In this paper, we study four classical tests of Schwarzschild space-time with Lorentzian distribution in non-commutative geometry. We performed detailed calculations of the first-order corrections induced by the non-commutative parameter on planetary orbital precession, light deflection, radar wave delay, and gravitational redshift. The study showed that the impact of the non-commutative parameter on timelike geodesics is significantly greater than its effect on null geodesics. Using precise experimental observations, the allowable range for the non-commutative parameter is ultimately constrained within $\Theta\leq 0.067579\leavevmode\nobreak\ \mathrm{m}^{2}$, which is given by Mercury’s orbital precession. This result aligns with the view that $\sqrt{\Theta}$ is on the order of the Planck length. Moreover, the constrained parameter range exceeds the Planck scale by a significant margin.
@@ I Introduction
Non-commutative geometry is a kind of quantum gravity theories noncomu1 ; noncomu2 ; noncomu3 ; noncomu4 ; noncomu5 ; noncomu6 ; noncomu7 . Unlike in normal quantum mechanics, non-commutative geometry posits that space-time coordinate operators themselves are non-commutative, as described by a relation $\left[\hat{x}^{\mu},\hat{x}^{\nu}\right]=\mathrm{i}\hat{\Theta}^{\mu\nu}$, with $\hat{\Theta}^{\mu\nu}$ being an anti-symmetric constant matrix. The parameter $\sqrt{\Theta}$ represents the minimum scale of space-time. While the precise value of $\Theta$ remains undetermined, $\sqrt{\Theta}$ is generally considered to be on the order of the Planck length commutator ; planck-length . There have been extensive research conducted to study the influences brought by non-commutative geometry in theories of gravity gr1 ; gr2 ; gr3 ; gr4 ; gr5 ; gr6 ; gr7 ; gr8 . An especially intriguing perspective is that, due to non-commutative effects, the mass distribution of a point particle, originally concentrated at a single point, is instead modified as a continuous distribution spread throughout the entire space. This offers a theoretical avenue for potentially addressing the black hole singularity problem. Two alternative forms of such matter distributions are commonly considered: the Lorentzian distribution $\rho=\frac{M}{\left(4\pi\Theta\right)^{\frac{3}{2}}}\exp\left(-\frac{r^{2}}{4\Theta}\right)$ lorentziandistribution and the Gaussian distribution $\rho=\frac{M\sqrt{\Theta}}{\pi^{\frac{3}{2}}\left(r^{2}+\pi\Theta\right)^{2}}$ gaussiandistribution1 ; gaussiandistribution2 ; gaussiandistribution3 . When $\Theta$ approaches zero, these two distributions both converge to a Dirac delta function $M\delta^{3}\left(x\right)$, leading to the standard Schwarzschild space-time. There have been many studies investigating the properties of black holes in the background of non-commutative geometry gr1 ; gr2 ; gr3 ; gr4 ; lorentziandistribution ; gaussiandistribution1 ; gaussiandistribution2 ; gaussiandistribution3 ; Gaussianthermo ; optics ; thermo1 ; thermo2 ; thermo3 ; thermo4 ; Horizon-scale ; gr9 ; gr10 ; gr11 ; gr12 ; gr13 . The effect of noncommutative geometry on the classical orbits of particles in a central force potential was investigated in centralpotential . Moreover, because the commutator of coordinate operators being a nonzero constant $\Theta$ matrix is incompatible with Lorentz symmetry noncomu1 , non-commutative geometry may offer valuable insights into the breaking of Lorentz covariance, which has also attracted much interest of researchers. lorentz1 ; lorentz2 ; lorentz2 ; lorentz3 ; lorentz4 ; lorentz5 ; lorentz6 .
@@
Since its inception, general relativity has undergone extensive experimental verification. Four of the most renowned tests include the precession of Mercury’s orbit, light deflection, gravitational redshift, and radar wave delay. Due to the gravitational effects, light will undergo a deflection when it passes near massive celestial bodies. It is noteworthy that the coordinate speed of light is not necessarily equal to the speed of light in a vacuum $c$. Combined with the fact that light bends in a gravitational field, this causes an observer to detect a time delay when receiving a radar wave emitted from their location and then reflected back. On the other hand, the variation in clock rates at different positions within a gravitational field results in a frequency shift of light waves, as they represent oscillations of the electromagnetic field. Additionally, because the time-like geodesics in Schwarzschild space-time does not satisfy the Binet equation in Newton’s mechanics, a correction at the scale of the Schwarzschild radius emerges. Consequently, the motion trajectory of a planet in Schwarzschild space-time deviates from a closed elliptical orbit, with its semi-major axis shifting periodically. Although $\sqrt{\Theta}$ is considered to be on the order of the Planck length $\ell_{p}(\simeq 1.616\times 10^{-35}\mathrm{m})$, it is still crucial to investigate the possible values of the non-commutative parameter. In BTZgravastar , the authors study the relationship between the noncommutative parameter and the cosmological constant by constructing a BTZ gravastar in non-commutative geometry, and obtain a relatively precise parameter constraint.
@@
In this work, we study the parameter constraints of Lorentzian distribution in non-commutative geometry using solar system tests. Investigating the constraints on this parameter provided by the classic experiments in the Solar System is of meaningful interest. Moreover, calculating the four classical tests of general relativity within the framework of non-commutative geometry is valuable for studying the impact of non-commutative geometry on the motion of time-like particles and light rays. This article is organized as follows. In Sect. II, we substitute the Lorentzian distribution into the Einstein equation to obtain the metric of Schwarzschild space-time in non-commutative geometry. In Sect. III, we conduct detailed calculations of four classical experiments in Schwarzschild space-time with the Lorentzian distribution, including planetary orbital precession, light deflection, radar wave delay, and gravitational redshift effects. By selecting more precise experimental values, we determine the range of the non-commutative parameter. Finally, we give our conclusion and outlook in Sect. IV.
@@ II Schwarzschild space-time of Lorentzian distribution in non-commutative geometry
where $M$ is the mass of the central celestial body and $\Theta$ is non-commutative parameter with dimension of $\left[\mathrm{L}^{2}\right]$. One could verify that
@@
It is evident that the metric asymptotically approaches the Schwarzschild geometry at infinity.
Specifically, when $r$ goes to zero,
@@ III Parameter constraints given by the solar system tests
Before studying the specific properties of this space-time, we first investigate the range of non-commutative parameter using the solar system test. Although the order of $\sqrt{\Theta}$ is considered to be at the scale of the Planck length $\ell_{p}$ commutator ; lp2 , there is currently a lack of a specific range for the non-commutative parameter. Therefore, it is worth studying the parameter constraints beforehand. In this section, the International System of Units (SI) is used. The metric in this case is
@@ III.1 Perihelion precession of planet
For a time-like particle, it line element $\mathrm{d}s^{2}\left(<0\right)$ is
@@
By introducing transformation $r=u^{-1}$, above function becomes
@@
Substituting the metric $f(r)$ then one could get the orbit equation of particle in non-commutative Schwarzschild space-time as follows
@@
The Binet equation, which works in Newton’s mechanics, is written as
@@
where $p=\frac{h^{2}}{GM}$ and $e$ is orbit’s eccentricity. Assuming the solution of Eq. 22 is $u=u_{0}+r_{\mathrm{S}}u_{1}$ ($r_{\mathrm{S}}=\frac{2GM}{c^{2}}$ is Schwarzschild radius), one could derive
@@
Considering that $a$ and $r_{\mathrm{S}}$ are small, one can keep above equation to first-order small quantity $\mathcal{O}\left(a,r_{\mathrm{S}}\right)$ and obtain
@@
Substituting $u_{1}=c_{1}+c_{2}\phi\sin\phi+c_{3}\cos 2\phi$, one could obtain three undetermined coefficients:
@@
Now we study the precession. In above expression, the perturbation of constant term will only cause a small deviation from original orbit $u_{0}$ and periodic term $\cos 2\phi$ doesn’t make a long-term correction of orbit. Only term $\phi\sin\phi$ plays a significant role as $\phi$ increases. So one could only consider this effective orbit equation
@@
where $R=\frac{p}{1-e^{2}}$ is semimajor axis of elliptic orbit, $T$ is period. Then Eq. 35 reads
@@
It is worth noticing that in above formula, the precession angle $\Delta_{\mathrm{c}}$ is measured in arcseconds ′′, the semimajor axis $R$ is measured in astronomical units AU (1AU=149597870700 m), the period $T$ is measured in sidereal years (1 yr=365.256 days), and the non-commutative geometry parameter $a$ is measured in meters m. For Mercury, its semimajor axis is $R=0.38709893$ AU, its orbital eccentricity is $e=0.20563069$ and its orbit period is $T=87.969$ days nasadata . And Mercury’s precession was measured as $\Delta_{\mathrm{c}}=\left(42.979\pm 0.003\right)^{\prime\prime}\mathrm{Century}^{-1}$ Mercury . It gives this following constraint
@@ III.2 Deflection of light
Now we continue to study the deflection of light. One could only replace parameter $s$ in Eqs. 16 and 17 with the affine parameter $\lambda$ to obtain the equation of motion for light
@@
Neglect the higher-order terms of $r_{\mathrm{S}}$, then $u_{1}$ should satisfy
@@
Without loss of generality, we set $u^{\prime}\left(\frac{\pi}{2}\right)=0$, which means photon reaches its perihelion when $\phi=\frac{\pi}{2}$. It gives the last undetermined coefficient $c_{2}=-\frac{3a\pi}{16b^{3}}$.
@@
For light just grazing the edge of the Sun, $M=M_{\odot}$ and $b=R_{\odot}$.
@@ III.3 Time delay of radar echo
[…]
@@
$1-\frac{f\left(r\right)}{f\left(r_{0}\right)}\frac{r_{0}^{2}}{r^{2}}=\left(1-\frac{r_{0}^{2}}{r^{2}}\right)\left(1+\frac{GM\left(ar+ar_{0}-2rr_{0}\right)}{c^{2}r^{2}\left(r+r_{0}\right)}\right)+\mathcal{O}\left(c^{-4}\right).$ (64)
Then one have
@@
Assuming a radar wave emitted from the Earth, travels through the Sun’s gravitational field, reaches a planet on the other side of the Sun, and then returns to the Earth along the same path. The total time of the radar echo is
@@ III.4 Gravitational redshift
Consider an atom emitting a beam of light, which propagates from the surface of the Sun to an observer on the Earth. Since the wavenumber $n=\nu\mathrm{d}\tau$ is a conserved quantity, where $\nu$ represents the frequency of the light and $\mathrm{d}\tau$ denotes the proper time interval, one could have
@@
where $\nu_{1}$ is the proper frequency of the light wave emitted by a stationary atom on the surface of the Sun, while $\nu_{2}$ is the frequency measured by a stationary observer on the Earth. $\mathrm{d}\tau_{1}$ and $\mathrm{d}\tau_{2}$ are the proper time intervals at the surface of the Sun and at the Earth, respectively, within the gravitational field of the Sun. With the relation $\mathrm{d}s^{2}=-c^{2}\mathrm{d}\tau^{2}=-c^{2}f\left(r\right)\mathrm{d}t^{2}$, one have
@@ III.5 A brief summary
From Eqs. 35, 55, 70 and 79, one can observe that non-commutative geometry exerts a weakening effect on gravitational interactions. Furthermore, the influence of non-commutative geometry on the geodesic motion of timelike particles is much more pronounced than its effect on the motion of photons. Tab. 1 is a summary of parameter constraints given by the solar tests. It is evident that the value of the non-commutative parameter inferred from solar system experiments should lie within the range of $\Theta\leq 0.067579\leavevmode\nobreak\ \mathrm{m}^{2}$. The scale of the Planck constant clearly falls within this range. In comparison, the range of $\Theta$ provided by the experiments appear to be excessively large. However, considering that the motions discussed in this section are located outside the solar surface, even for the result with the maximum width among the four datasets, there remains $\sqrt{1.6358\times 10^{13}}\leavevmode\nobreak\ \mathrm{m}/R_{\odot}=5.8\times 10^{-3}\ll 1$. Therefore, employing Eq. 6 for computations is deemed sufficiently safe. And, on astronomical scales, these constraints already represent relatively precise limits.

### v4

@@ Abstract
In this paper, we study four classical tests of Schwarzschild space-time with Lorentzian distribution in non-commutative geometry. We performed detailed calculations of the first-order corrections induced by the non-commutative parameter on planetary orbital precession, light deflection, radar wave delay, and gravitational redshift. The study showed that the impact of the non-commutative parameter on the time-like geodesics is significantly greater than its effect on the null geodesics. By using a series of precise experimental observations, the allowable range for the non-commutative parameter is ultimately constrained within $\Theta\leq 0.067579\leavevmode\nobreak\ \mathrm{m}^{2}$, which is given by Mercury’s orbital precession. This result aligns with the view that $\sqrt{\Theta}$ is of the order of the Planck length. Moreover, this constrained parameter range exceeds the Planck scale by a significant margin.
@@ I Introduction
Modified gravity theory is a classic and widely studied research direction in the field of general relativity (GR) mg1 ; mg2 ; mg3 ; mg4 ; mg5 ; mg6 ; mg7 ; mg8 ; mg9 ; mg10 ; mg11 . Like quantum gravity qg1 ; qg2 ; qg3 ; qg4 ; qg5 ; qg6 ; qg7 ; qg8 , non-commutative geometry is also a theory of gravitational quantization noncomu1 ; noncomu2 ; noncomu3 ; noncomu4 ; noncomu5 ; noncomu6 ; noncomu7 . Unlike in traditional quantum mechanics, non-commutative geometry posits that space-time coordinate operators themselves are non-commutative, as described by a relation $\left[\hat{x}^{\mu},\hat{x}^{\nu}\right]=\mathrm{i}\hat{\Theta}^{\mu\nu}$, where $\hat{x}^{\mu}$ is space-time coordinate operator and $\hat{\Theta}^{\mu\nu}$ is an anti-symmetric constant matrix. The parameter $\sqrt{\Theta}$ represents the minimum scale of space-time. While the precise value of $\Theta$ remains undetermined, $\sqrt{\Theta}$ is generally considered to be of the order of the Planck length commutator ; planck-length . There have been extensive research conducted to study the influences brought by non-commutative geometry in theories of gravity gr1 ; gr2 ; gr3 ; gr4 ; gr5 ; gr6 ; gr7 ; gr8 . An especially intriguing perspective is that, due to non-commutative effects, the mass distribution of a point particle, originally concentrated at a single point, is instead modified as a continuous distribution spreading throughout the entire space. This offers a theoretical avenue for potentially addressing the black hole singularity problem. Two alternative forms of such matter distributions are commonly considered: the Lorentzian distribution $\rho=\frac{M}{\left(4\pi\Theta\right)^{\frac{3}{2}}}\exp\left(-\frac{r^{2}}{4\Theta}\right)$ lorentziandistribution and the Gaussian distribution $\rho=\frac{M\sqrt{\Theta}}{\pi^{\frac{3}{2}}\left(r^{2}+\pi\Theta\right)^{2}}$ gaussiandistribution1 ; gaussiandistribution2 ; gaussiandistribution3 . When $\Theta$ approaches zero, these two distributions both converge to a Dirac delta function $M\delta^{3}\left(x\right)$, leading to a standard Schwarzschild space-time. There have been many studies investigating the properties of black holes in the background of non-commutative geometry gr1 ; gr2 ; gr3 ; gr4 ; lorentziandistribution ; gaussiandistribution1 ; gaussiandistribution2 ; gaussiandistribution3 ; Gaussianthermo ; optics ; thermo1 ; thermo2 ; thermo3 ; thermo4 ; Horizon-scale ; gr9 ; gr10 ; gr11 ; gr12 ; gr13 . The effect of noncommutative geometry on the classical orbits of particles in a central force potential was investigated in centralpotential . Moreover, because the commutator of coordinate operators being a nonzero constant $\Theta$ matrix is incompatible with Lorentz symmetry noncomu1 , non-commutative geometry may offer valuable insights into the breaking of Lorentz covariance, which has also attracted much interest of researchers. lorentz1 ; lorentz2 ; lorentz2 ; lorentz3 ; lorentz4 ; lorentz5 ; lorentz6 .
@@
GR has undergone extensive experimental verification. Four of the most renowned tests include the precession of Mercury’s orbit, light deflection, gravitational redshift, and radar wave delay. Due to the gravitational effect, the light will undergo a deflection when it passes near a massive celestial body. Additionally, considering the effect caused by the gravitational slowing down of light, the motion of light in a gravitational field will result in a time delay. And, the time delay can happen even when light travels in a radial direction. On the other hand, the variation in clock rates at different positions within a gravitational field results in a frequency shift of light waves, as they represent oscillations of the electromagnetic field. Additionally, because the time-like geodesics in Schwarzschild space-time does not satisfy the Binet equation in Newton’s mechanics, a correction at the scale of the Schwarzschild radius emerges. Consequently, the trajectory of a planet in Schwarzschild space-time is no longer a closed elliptical orbit, with its semi-major axis shifting periodically. Although $\sqrt{\Theta}$ is considered to be on the order of the Planck length $\ell_{p}(\simeq 1.616\times 10^{-35}\mathrm{m})$, it is still crucial to investigate the limit of the non-commutative parameter. In BTZgravastar , the authors study the relationship between the noncommutative parameter and the cosmological constant by constructing a BTZ gravastar in non-commutative geometry, and obtain a relatively precise parameter constraint.
@@
In this work, we study the parameter constraints of Lorentzian distribution in non-commutative geometry using solar system tests. Investigating the constraints on this parameter provided by the classic experiments in the Solar System is of meaningful interest. Moreover, calculating the four classical tests of GR within the framework of non-commutative geometry is valuable for studying the impact of non-commutative geometry on the motion of time-like particles and light rays. This article is organized as follows. In Sect. II, we substitute the Lorentzian distribution into the Einstein equation to obtain the metric of Schwarzschild space-time in non-commutative geometry. In Sect. III, we conduct detailed calculations of the four classical experiments in Schwarzschild space-time with the Lorentzian distribution, including planetary orbital precession, light deflection, radar wave delay, and gravitational redshift effects. With the help of a series of precise observational data, we determine a range of the non-commutative parameter. Finally, we give our conclusion and outlook in Sect. IV.
@@ II Schwarzschild space-time of Lorentzian distribution in non-commutative geometry
where $M$ is the mass of the central celestial body and $\Theta$ is the non-commutative parameter with dimension of $\left[\mathrm{L}^{2}\right]$. One could verify that
@@
[…]
Specifically, when $r$ goes to zero, one have
@@ III Parameter constraints given by the solar system tests
In this section, we separately calculate the perihelion precession of Mercury, light deflection, time delay of light, and gravitational redshift in Schwarzschild space-time with the Lorentzian distribution. Additionally, we choose some of the more precise observational data to derive an acceptable range for the non-commutative parameter. In this section, the International System of Units (SI) is used. The metric in this case is
@@ III.1 Perihelion precession of planet
For a time-like particle, its line element $\mathrm{d}s^{2}\left(<0\right)$ is
@@
By introducing transformation $r=u^{-1}$, the above function becomes
@@
By substituting the metric $f(r)$, one could get the orbit equation of particles in non-commutative Schwarzschild space-time as follows
@@
Please notice that the Binet equation, which works in Newton’s mechanics, is
@@
where $p=\frac{h^{2}}{GM}$ and $e$ is the orbit’s eccentricity. Assuming the solution of Eq. 22 is $u=u_{0}+r_{\mathrm{S}}u_{1}$ ($r_{\mathrm{S}}=\frac{2GM}{c^{2}}$ is Schwarzschild radius), one could derive
@@
Considering that $a$ and $r_{\mathrm{S}}$ are small, one can keep above equation to the first-order small quantity $\mathcal{O}\left(a,r_{\mathrm{S}}\right)$ and obtain
@@
Substituting $u_{1}=c_{1}+c_{2}\phi\sin\phi+c_{3}\cos 2\phi$, one could obtain the undetermined coefficients:
@@
Now we study the precession. In above expression, the perturbation of constant term will only cause a small deviation from the original orbit $u_{0}$ and the periodic term $\cos 2\phi$ doesn’t make a long-term correction of the orbit. Only the term $\phi\sin\phi$ plays a significant role as $\phi$ increases. So one could only consider this effective orbit equation
@@
where $R=\frac{p}{1-e^{2}}$ is semi-major axis of elliptic orbit, $T$ is period. Then Eq. 35 reads
@@
It is worth noticing that in above formula, the precession angle $\Delta_{\mathrm{c}}$ is measured in arcseconds ′′, the semi-major axis $R$ is measured in astronomical units AU (1AU=149597870700 m), the period $T$ is measured in sidereal years (1 yr=365.256 days), and the non-commutative parameter $a$ is measured in meters m. For Mercury, its semi-major axis, orbital eccentricity, and orbit period are $R=0.38709893$ AU, $e=0.20563069$ annd $T=87.969$ days respectively nasadata . Regarding the experimental tests of gravitational theory within the solar system, it is worth mentioning E.V. Pitjeva, A. Fienga, and L. Iorio. In recent years, each of their teams has conducted a considerable amount of valuable work on this subject exp1 ; exp2 ; exp3 ; exp4 ; exp5 ; exp6 ; exp7 ; exp8 ; exp9 ; exp10 ; exp11 . In this paper, we just take the data from Mercury as an example to estimate the upper limit of $\Theta$. Mercury’s precession was reported as $\Delta_{\mathrm{c}}=\left(42.979\pm 0.003\right)^{\prime\prime}\mathrm{Century}^{-1}$ Mercury . It gives this following constraint
@@ III.2 Deflection of light
Now we continue to study the deflection of light. One could only replace the parameter $s$ in Eqs. 16 and 17 with the affine parameter $\lambda$ to obtain the equation of motion for light
@@
Neglecting the higher-order terms of $r_{\mathrm{S}}$, then $u_{1}$ should satisfy
@@
Without loss of generality, we set $u^{\prime}\left(\frac{\pi}{2}\right)=0$, which means photon reaches its perihelion at $\phi=\frac{\pi}{2}$. It gives the last undetermined coefficient $c_{2}=-\frac{3a\pi}{16b^{3}}$.
@@
For light just grazing the edge of the Sun, one have $M=M_{\odot}$ and $b=R_{\odot}$.
@@ III.3 Time delay of radar echo
To study the time delay of light in gravitational field, we use the method illustrated in RNtime . In this reference, the post-post-Newtonian amendment is applied to calculate the time delay in RN space-time. The inclusion of $\mathcal{O}\left(c^{-4}\right)$ is necessary for calculating the time delay in RN space-time because the electric charge $Q$ only becomes significant at this magnitude. In this paper, the post-Newtonian approximation is chosen because the non-commutative parameter $a$ already becomes obvious when considering the magnitude of $\mathcal{O}\left(c^{-2}\right)$.
@@
$1-\frac{f\left(r\right)}{f\left(r_{0}\right)}\frac{r_{0}^{2}}{r^{2}}=\left(1-\frac{r_{0}^{2}}{r^{2}}\right)\left(1+\frac{GM\left(ar+ar_{0}-2rr_{0}\right)}{c^{2}r^{2}\left(r+r_{0}\right)}\right)+\mathcal{O}\left(c^{-4}\right),$ (64)
one have
@@
Assuming that a radar wave emits from the Earth, traveling through the Sun’s gravitational field, then reaches a planet on the other side of the Sun, and finally returns back to the Earth along the same path. The total time of the radar echo is
@@ III.4 Gravitational redshift
Considering that an atom emitting a beam of light, which propagates from the surface of the Sun to an observer on the Earth. Since the wavenumber $n=\nu\mathrm{d}\tau$ is a conserved quantity, where $\nu$ represents the frequency of the light and $\mathrm{d}\tau$ denotes the proper time interval, one could have
@@
where $\nu_{1}$ is the proper frequency of the light wave emitted by a stationary atom on the surface of the Sun, while $\nu_{2}$ is the frequency measured by a stationary observer on the Earth. $\mathrm{d}\tau_{1}$ and $\mathrm{d}\tau_{2}$ are the proper time intervals at the surface of the Sun and at the Earth, respectively. With the relation $\mathrm{d}s^{2}=-c^{2}\mathrm{d}\tau^{2}=-c^{2}f\left(r\right)\mathrm{d}t^{2}$, one have
@@ III.5 A brief summary
From Eqs. 35, 55, 70 and 79, one can observe that non-commutative geometry exerts a weakening effect on the gravitational interactions. Furthermore, the influence of non-commutative geometry on the geodesic motion of time-like particles is much more pronounced than its effect on the motion of photons. Tab. 1 is a summary of parameter constraints given by the solar tests. It is evident that the value of the non-commutative parameter inferred from solar system experiments should lie within the range of $\Theta\leq 0.067579\leavevmode\nobreak\ \mathrm{m}^{2}$. The scale of the Planck length clearly falls within this range. In comparison, the range of $\Theta$ provided by the experiments appear to be excessively large. However, considering that the motions discussed in this section are located outside the solar surface, even for the result with the maximum width among the four datasets, there remains $\sqrt{1.6358\times 10^{13}}\leavevmode\nobreak\ \mathrm{m}/R_{\odot}=5.8\times 10^{-3}\ll 1$. Therefore, employing Eq. 6 for computations is deemed sufficiently safe. And, on astronomical scales, these constraints already represent relatively precise limits.